If in the process oof rubbing the lenses of the eyeglasses, 6.28 x 10^10 electrons were transferred,


a) what is the charge of the lenses and the nylon cloth?


b) what is the change in their masses?

1
Expert's answer
2022-02-23T08:31:16-0500

The charge of the lenses (glass) is positive


"q_1=eN\\
=1.6\times 10^{-19}\:\rm C\times 6.28\times 10^{10}=1.00\times 10^{-8}\:\rm C"

The charge of the nylon cloth is negative


"q_2=-eN\\
=-1.6\times 10^{-19}\:\rm C\times 6.28\times 10^{10}=-1.00\times 10^{-8}\:\rm C"


The change in mass

"\Delta m=Nm_e\\
=6.28*10^{10}*9.1*10^{-31}\:\rm kg=5.71*10^{-20}\: kg"

The lenses lose, the nylon cloth gains.